OpenStudy (anonymous):

Find the greatest possible product xy given that x and y are both positive integers and x + y = 62 @Calculus1

OpenStudy (anonymous):

31 * 31 = 961

OpenStudy (anonymous):

The least product will be 60 * 1 = 60

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Welcome

OpenStudy (radar):

60 * 1 60 + 1 does not equal 62.

OpenStudy (radar):

but 31 is the correct answer.

OpenStudy (radar):

do you know how pratu043 solved this?

OpenStudy (anonymous):

yes, with substitution to find x and y using x+y=62 and the derivative of the area of a square. Then multiplied x and y together.

OpenStudy (radar):

Yes, good luck with these.

OpenStudy (anonymous):

thank you